Question 2 Partially correct Mark 1.00 out of 6.00 Remove flag Edit question Compute and Interpret Measures for DuPont DisaggCould Someone help me out with this? I think its write but It says its wrong
Answer #1

a)  ROE:    Net Income / Average common equity

= 5349 / [9848+11622] / 2

= 49.83%

b) Dupont's Profit Margin ; Net income / sales

= 5,363 / 32,765

= 16.37

Dupont's Financial Leverage :

Average Total assets / Average shareholders equity

   = [ (36,500+37,987) / 2 ] / [ (9,796+11,563) / 2 ]

= 37,243.50 / 10,679.50

= 34.875

  

c) ROA : Net Income / Average Total Assets

= 5349 / [ (36,500+37,987) / 2 ]

= 5349 / 37,243.50

= 14.36

d) Adjusted ROA : [ Net Income - Pretax Interest Expense - Tax ] / Average Total Assets

Net income =   5349

( - ) Pre tax Interest = (207)

_____________

5,132

( - ) Tax @ 22 % (1,129.04)

_______________

4,002.96

Adjusted ROA = 4,002.96 / [ (36,500+37,987) / 2 ]

= 4,002.96 /  37,243.50

= 10.75